### v2

# Cybersecurity Defenses: Exploration of CVE Types through Attack Descriptions

Abstract
Vulnerabilities in software security can remain undiscovered even after being exploited. Linking attacks to vulnerabilities helps experts identify and respond promptly to the incident. This paper introduces VULDAT, a classification tool using a sentence transformer MPNET to identify system vulnerabilities from attack descriptions. Our model was applied to 100 attack techniques from the ATT&CK repository and 685 issues from the CVE repository. Then, we compare the performance of VULDAT against the other eight state-of-the-art classifiers based on sentence transformers. Our findings indicate that our model achieves the best performance with F1 score of 0.85, Precision of 0.86, and Recall of 0.83. Furthermore, we found 56% of CVE reports vulnerabilities associated with an attack were identified by VULDAT, and 61% of identified vulnerabilities were in the CVE repository.
Index Terms:
MITRE, CAPEC, CVE, Transformer models, Pretrained language models
I Introduction
Cyberattacks are expected to cost organizations \$3 trillion in 2015, \$6 trillion in 2021, and more than \$10.5 trillion annually by 2025 [1]. Moreover, an attacker targets the organization’s system over a thousand times a week on average [2]. However, Cyber Threat Intelligence (CTI) can be utilized by cybersecurity practitioners [3] [4], which refers to the information that practitioners can use to protect themselves proactively against cyberattacks. The MITRE Corporation, for instance, has created a collection of resources, such as Tactics Techniques and Common Knowledge repository (ATT&CK)  [5], Common Attack Pattern Enumeration and Classification (CAPEC) [6], Common Weakness Enumeration (CWE) [7], and Common Vulnerabilities and Exposures (CVE) [8]. ATT&CK contains descriptions of attacks as Tactics, Techniques, and Procedures (TTP). CAPEC is called an Attack Pattern, which is “the common approach and attributes related to exploiting a weakness in a software, firmware, hardware, or service component” [6]. Additionally, CAPEC is structured as a list of information about the attack, including techniques for the attack, potential outcomes, and mitigations [9]. CWE is a community-developed collection of typical weaknesses in software, coding errors, and security flaws. CVE is a standardized dictionary of common terms for publicly known cybersecurity vulnerabilities. Its initiative aims to enhance the efficiency of identifying, finding, and fixing software vulnerabilities by providing a unified naming system [10] [11]. The CVE entries are widely used, but CVE issues often lack information regarding mitigation techniques or existing defense strategies that could effectively address a specific vulnerability [12, 13]. Linking ATT&CK to CVE issues will help the cyber community, as these two repositories are currently separated. However, linking 625 Attack Techniques manually with 295,604 CVE issues  [8] is a non-trivial task requiring automated models to link attack information with vulnerabilities and weaknesses. This study aims to develop an approach that links the Attack Techniques (AT) in the ATT&CK repository to reports in the CVE repository. To achieve this goal, we developed a tool named Automated Vulnerability Detection From Attack Technique Text (VULDAT), that leverages the sentence transformer (MPNET) [14] that connects the textual description of an AT to the textual description of CVE reports. Thus, we aim to answer the following questions:
  • RQ1: What is the performance of sentence transformer models to detect software vulnerabilities from attack text?
  • RQ2: How many CVE issues VULDAT correctly detect from an attack text?
The contributions of this work are as follows:
  • The proof-of-concept VULDAT supports cybersecurity by providing an automated way of detecting software vulnerabilities from AT text;
  • A novel annotated mapping dataset  [15] explicitly linking ATT&CK with vulnerabilities found in MITRE repositories;
  • An analysis of sentence transformer models, highlighting their comparative features and performance metrics;
The structure of this paper is as follows. Section II illustrates our proposed methodology for linking attacks to CVE. Section III summarizes the results of our work. Section V outlines the limitations of our work. Finally, section VI presents our conclusions and further research.
Figure 1: Methodology Overview
II Study Design and Methodology
In this section, we describe our research methodology for vulnerability detection. Fig. 1 illustrates the implemented pipeline.
II-A Data Collection
Table I and Fig. 2 provide an overview of the Mapping dataset $\mathcal{M}$ used in this paper and summarize the number of collected items from MITRE repositories, ATT&CK  [5], CAPEC  [6], CWE  [7], and CVE  [8]. ATT&CK and CAPEC reports are connected through the technique ID, CAPEC and CWE reports are connected through the CWE-ID, and CWE and CVE reports are connected through the CVE-ID. The dataset provides extensive vulnerability information for effective cybersecurity vulnerability management and prioritization.
TABLE I: ATT&CK linked and not linked to vulnerability reports
Linked Not linked Total
Attack Patterns 177 382 559
CWE reports 117 817 934
CVE reports 685 294919 295604
For instance, we found that only 100 attack techniques/sub-techniques are linked to 88 CWE reports and the same CWE reports are connected to 685 CVE reports. Meanwhile, a substantial number of 293,589 CVE issues, 370 CWEs, 282 CAPEC, and 230 Techniques are isolated and are not linked to any other information, as shown in Fig. 2.
Figure 2: MITRE repositories and their connections
II-B Pre-Processing
Fig. 3 illustrates the steps of pre-processing. The pre-processing phase in our experiment involves two parallel steps: (1) Our initial pre-processing (partial pre-processing) experiment includes the same steps as the pretrained model: tokenization, stemming, lemmatization, stop word removal, and punctuation handling. (2) Additional pre-processing that removes citations, URLs, extra spaces, and non-alphanumeric characters (full pre-processing). Tokenization involves splitting the text into tokens or single words. This process allows us to handle the text more effectively, making extracting meaningful insights and patterns from the data easier. Stemming and Lemmatization are the process of breaking down a word into its stem or root form. This process aims to unify variations of words that share the same core meaning, reducing the complexity of the data and improving our analysis. Moreover, stop words are removed from the textual descriptions of the attack information, as they are often considered irrelevant or noisy for natural language processing tasks. We remove citations, URLs, extra spaces, and non-alphanumeric characters from AT texts to guarantee a cleaner dataset.
Figure 3: Text pre-processing
II-C Designing VULDAT Architecture
In this paper, we use state-of-the-art pre-trained sentence transformer models [16] as summarized in Table II. These models are based on the Distilbert  [17], Bert [18], Roberta [19], and Mpnet [20] architectures. These include multi-qa-MiniLM-L6-cos-v1 and all-MiniLM-L12-v2, which utilize the MiniLM architecture for multilingual and cross-lingual tasks.
TABLE II: Pretrained Model Summary
Model Embedding Dimension Model Size
multi-qa-MiniLM-L6-cos-v1 384 80   MB
multi-qa-distilbert-cos-v1 768 250 MB
all-MiniLM-L12-v2 384 120 MB
all-distilroberta-v1 768 290 MB
multi-qa-mpnet-base-dot-v1 768 420 MB
all-MiniLM-L6-v2 384 80   MB
paraphrase-multilingual-MiniLM-L12-v2 384 420 MB
all-mpnet-base-v2 768 420 MB
paraphrase-MiniLM-L6-v2 384 61   MB
Furthermore, multi-qa-distilbert-cos-v1 and all-distilroberta-v1, which are based on the DistilBERT and DistilRoBERTa architectures, respectively. Using the MPNet architecture, the multi-qa-mpnet-base-dot-v1 and all-mpnet-base-v2. Lastly, the MiniLM architecture, all-MiniLM-L6-v2, paraphrase-multilingual-MiniLM-L12-v2, and paraphrase-MiniLM-L6-v2 models, are designed for use in multilingual paraphrasing tasks. The sentence transformer models generate embeddings for the AT and the CVE descriptions. Then, we calculate a cosine similarity score between these vectors, which measures how closely the CVE descriptions match the AT text. Thus, the similarity score helps identify and assign the most relevant CVEs to each AT.
II-D Compute Similarity Score
We apply our model to associate an AT with CVE report vectors by leveraging their cosine similarity scores. The output CVE vectors are ranked based on their similarity to the AT vector. The top N CVE reports with a similarity score greater than 60% are then associated with the AT vector. This set of CVE reports, named the Detection List ($\mathcal{L}$), represents the most relevant vulnerabilities linked to the AT.
II-E Performance Evaluation
[…]
TABLE III: Positives and Negatives. $\mathcal{A}$ = Set of all Attacks, C = Set of all CVE issues
Type Description
Positives $\{a\in\mathcal{A}\,:\exists\,{c}\in\mathcal{M}_{a}\}$
Negatives $\{a\in\mathcal{A}\,:\nexists\,{c}\in\mathcal{M}_{a}\}$
True Positive $\{a\in\mathcal{A}\,:\left|\mathcal{L}_{a}\cap\mathcal{M}_{a}\right|>0\}$
False Positive $\{a\in\mathcal{A}\,:\left(\left|\mathcal{L}_{a}\right|>0\,\land\,\left|\mathcal{M}_{a}\right|=0\right)\}$
False Negative $\{a\in\mathcal{A}\,:\left(\left|\mathcal{L}_{a}\right|=0\,\land\,\left|\mathcal{M}_{a}\right|>0\right)\}$
True Negative $\{a\in\mathcal{A}\,:\left(\left|\mathcal{L}_{a}\right|=0\,\land\,\left|\mathcal{M}_{a}\right|=0\right)\}$
Positives is when there is an explicit link between an AT text and a CVE, which means when we have $\mathcal{L}$ is not empty, while a negatives is when $\mathcal{L}$ is empty which means no explicit link exists.
III Results
This section presents the outcomes for each of the RQs.
RQ1: What is the performance of sentence transformer models to detect software vulnerabilities from attack text?
To answer RQ1, we evaluated the performance of the sentence transformer models in detecting CVE reports based on the techniques information from the ATT&CK repository under different pre-processing conditions (see Section II-B). Table  IV illustrates the precision, recall, and F1. Among the models evaluated, the multi-qa-mpnet-base-dot-v1 model achieved the best performance metrics under both partial and full pre-processing. With partial pre-processing, it demonstrated a precision of 0.73, a recall of 0.87, and an F1 score of 0.79. Furthermore, under full pre-processing, the model demonstrated significant improvements across all metrics, with an F1 score of 0.85, precision of 0.86, and recall of 0.83. However, the multi-qa-distilbert-cos-v1 model performed relatively poorly in comparison under either partial pre-processing with an F1 score of 0.32 or full
TABLE IV: Performance metrics of sentence transformer Models
Partial pre-processing Full pre-processing
Model Precision Recall F1 Precision Recall F1
multi-qa-MiniLM-L6-cos-v1 1.00 0.32 0.48 1.00 0.26 0.42
multi-qa-distilbert-cos-v1 1.00 0.19 0.32 0.88 0.13 0.22
all-MiniLM-L12-v2 0.87 0.62 0.72 0.91 0.54 0.68
all-distilroberta-v1 0.63 0.60 0.62 0.69 0.74 0.71
multi-qa-mpnet-base-dot-v1 0.73 0.87 0.79 0.86 0.83 0.85
all-MiniLM-L6-v2 0.95 0.38 0.54 0.74 0.26 0.38
paraphrase-multilingual-MiniLM-L12-v2 0.67 0.85 0.75 0.65 0.93 0.76
all-mpnet-base-v2 0.75 0.77 0.76 0.81 0.80 0.81
paraphrase-MiniLM-L6-v2 0.68 0.46 0.55 0.80 0.61 0.69
pre-processing with an F1 score of 0.22. In our case, the size of the model is associated with the number of its parameters and may have an impact on the performance of the models, as shown in Table II. A larger size allows the model to capture more complex patterns. For example, the multi-qa-mpnet-base-dot-v1 model has more parameters with a size of 420 MB, which allows it to work more efficiently. On the other hand, smaller models like the multi-qa-distilbert-cos-v1, which is 250 MB in size, have fewer parameters, which could lower performance. Thus, the F1 scores are 0.79 and 0.32 for partial pre-processing and 0.85 and 0.22 for full pre-processing.
RQ2: How many CVE issues VULDAT correctly detect from an attack text? To answer RQ2, we compute the Jaccard Similarity, Mapping Accuracy, and Detection Accuracy for each AT text. The Jaccard Similarity evaluates the similarity of two sets: the detected set ($\mathcal{L}_{a}$) by VULDAT and actual sets ($\mathcal{M}_{a}$). It measures the ratio of the intersection of these sets to their union, $\mathrm{Jaccard\,Similarity}=\frac{{|\mathcal{L}_{a}\cap\mathcal{M}_{a}|}}{|\mathcal{L}_{a}\cup\mathcal{M}_{a}|}$. Mapping Accuracy is the ratio of the intersection of the detected set ($\mathcal{L}_{a}$) and the actual set ($\mathcal{M}_{a}$) to the size of the actual set ($\mathcal{M}_{a}$), $\mathrm{Mapping\,Accuracy}=\frac{{|\mathcal{L}_{a}\cap\mathcal{M}_{a}|}}{{|\mathcal{M}_{a}|}}$. Detection Accuracy is the ratio of the intersection of the detected set ($\mathcal{L}_{a}$) and the actual set ($\mathcal{M}_{a}$) to the size of the detected set ($\mathcal{L}_{a}$), $\mathrm{Detection\,Accuracy}=\frac{{|\mathcal{L}_{a}\cap\mathcal{M}_{a}|}}{{|\mathcal{L}_{a}|}}$. A higher Detection Accuracy implies more accurate identifications by VULDAT.
Figure 4: VULDAT performance
Fig. 4 reports the distribution of these measures across all AT texts, showing a Mapping Accuracy of 0.56 and VULDAT’s reported Detection Accuracy of 0.61. Moreover, VULDAT reports a Jaccard Similarity of 0.40 on average, with a range spanning from 0.05 to 0.82. Whiskers extend further below Q1 and above Q3, indicating a slight skewness in the data distribution. Notably, most of the similarity scores are higher than 0.50, indicating that many sample comparisons demonstrate relatively high similarity. For instance, the T1539 attack achieves the highest Jaccard Similarity score of 0.82. Specifically, it consists of 150 CVE issues in $\mathcal{L}$, 125 CVE issues in $\mathcal{M}$, with 124 CVE issues in $\mathcal{L}_{a}\cap\mathcal{M}_{a}$, and 151 CVE issues at $\mathcal{L}_{a}\cup\mathcal{M}_{a}$. Additionally, a lower Jaccard Similarity value suggests limited agreement between the elements of the real and detected sets. Specifically, T1003 exhibits a Jaccard Similarity of 0.05. Upon manual evaluation of this attack by examining which CVE related to the AT text, VULDAT discovered the size of $\mathcal{L}_{T1003}$ is 37, and we found 22 of these links do not exist in the MITRE repositories. Thus, it is important to note that many detected CVE links may indeed be linked to the AT. However, their link is missing from the MITRE repositories. Our future plan to validate all of these $\mathcal{L}_{a}-\mathcal{M}_{a}$ links are indeed connected.
IV Related Work
The Cve2att&ck  [21] model uses BERT-based [18] language models and classic models for machine learning to automatically link CVE reports with techniques that are able to link CVE issues with the MITRE ATT&CK Enterprise Matrix techniques by utilizing the textual description present in CVE metadata. The CVE Transformer (CVET)  [22] model uses RoBERTa [19] to link a CVE to ten tactics from ATT&CK Enterprise Matrix. Kuppa et al. [23] introduced the Multi-Head Joint Embedding Neural Network model to automatically map CVE issues to ATT&CK techniques. The suggested representation finds particular regular expressions in the current threat reports and then compares the ATT&CK technique vectors with the text description found in the CVE metadata by measuring their similarity using the cosine distance. SMET  [24] is an automatic mapping tool that matches CVE entries to ATT&CK techniques using textual similarity. It uses the ATT&CK BERT model, trained on the SIAMESE network, to determine the semantic similarity between attack actions.
V Threats to Validity
In this section, we discuss the limitations of our research.
Construct validity. We determine the output of VULDAT by experimentally setting a 60% threshold. We plan to perform a sensitivity analysis on the threshold to substantiate our choice.
Internal validity. We exploit the description of the attack techniques. Other attack information included in the MITRE repository could eventually provide better performance results.
External validity. The link between attacks and vulnerabilities is based on what the MITRE repositories explicitly indicate. These links have not been updated with a stringent schedule. Thus, our dataset may not contain the very last pieces of information about an attack and its related vulnerabilities.
VI Conclusion
This paper introduces VULDAT, an approach that uses MPNET to detect CVE issues from textual descriptions of attack behaviors. It utilizes data from MITRE ATT&CK and CVE repositories. VULDAT’s performance was compared with nine state-of-the-art models and evaluated using three accuracy measures: Mapping, Detection, and Jaccard Similarity. VULDAT achieved a Mapping Accuracy of 0.56, a Detection Accuracy of 0.61, a Jaccard Similarity of 0.40, an F1 score of 0.85, a Precision of 0.86, and a Recall of 0.83. Manual inspection of its output confirmed some identified vulnerabilities were associated with attacks in the CVE repository. We plan to investigate other models (e.g., sequence-to-sequence models) for detecting CVE issues from attack text. Furthermore, we will perform a large manual inspection of VULDAT output to recommend missing links between attacks and vulnerabilities to the MITRE board.
VII Acknowledgements
The first author thanks the CSLab at the Free University of Bozen-Bolzano for supporting this work under project no. EFRE1039 in the EFRE-FESR 2021-2027 program.